OpenStudy (bbb911):

Solve. Round results to the nearest thousandth. 5x^2 - 14x + 14 = 0

OpenStudy (campbell_st):

well the question will have complex roots.... the discriminant is 14^2 - 4 x 5 x 14 = -84

OpenStudy (bbb911):

that what i have but the answers are different

OpenStudy (campbell_st):

so the solutions using the general quadratic formula are \[x = \frac{14 \pm \sqrt{-84}}{2 \times 5}\] which becomes \[x = \frac{7 \pm i \sqrt{21}}{5}\] hope that helps

OpenStudy (campbell_st):

well if it was - 14 then you will get real solutions...

OpenStudy (bbb911):

ok lemme show you the answers cos im confused

OpenStudy (bbb911):

a. 2.317, 0.483 b. 0.782, -3.582 c. 3.582, -0.782 d. -0.483, -2.317

OpenStudy (campbell_st):

well I got answer b when I used the equation \[5x^2 - 14x - 14 = 0\]

OpenStudy (campbell_st):

you get \[x = \frac{14 \pm \sqrt{476}}{10}\]

OpenStudy (campbell_st):

well I got the 1st part of b the 2nd option I had was a postive... not negative... same value...

OpenStudy (bbb911):

oh ok sounds good ill just put that as the answer thank you :D

OpenStudy (bbb911):

i was wonder if you can help me on 3 more questions :)

OpenStudy (campbell_st):

I think the equation should read \[5x^2 + 14x - 14 = 0\] if you use that solution I think you get option b this is definitely the equation that gives b as the answer looks like a typo

OpenStudy (campbell_st):

if you post them I'll look at them

OpenStudy (bbb911):

oh my goodness that is what i was thinking about it was a typo so i wasn't going crazy after all